compare the series formula and parallel formula in RESISTANCE, VOLTAGE, and CURRENT.


SERIES FORMULA

current -

voltage -

resistance-


PARALLEL FORMULA

current -

voltage-

resistance


1
Expert's answer
2021-05-24T09:00:53-0400

SERIES FORMULA

Let there be n resistance connected in series

Equivalent Resistance : "nR"

Current : "I"

Voltage : "nIR"


PARALLEL FORMULA

Let there be n resistance connected in parallel

Equivalent resistance : "\bigg(\dfrac{1}{R}+\dfrac{1}{R}+\dfrac{1}{R}+\dfrac{1}{R}+\dfrac{1}{R}.........+\dfrac{1}{R}\bigg)^{-1}"

"=\dfrac{R}{n}"

Current in each branch "= I_1"

Total current, "I" "=nI_1"

Voltage : "\dfrac{I_1R}{n}"